# The Origin of Quantum Mechanical Statistics: Some Insights from the Research on Human Language

### v1

@@ 3 Indistinguishability versus statistical independence
We intend in this section to establish a connection between distinguishability of identical entities and statistical independence of their relevant states. We will first consider the typical situation that occurs in physics (Gorroochurn,, 2018). Then, we will see that a completely analogous connection can be made in the case of human cognition, more specifically, language (Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ).
Traditional derivations of both classical, Maxwell–Boltzmann, and quantum, Bose–Einstein and Fermi–Dirac, distributions employ a method of distributing particles of the same type, which play the role of identical physical entities, across baskets, which play the role of micro-states (see, e.g., (Huang,, 1987; Darrigol,, 1991)). Let us consider a simple situation, namely, the distribution of two physical entities, labelled $1$ and $2$, in two micro-states, labelled $p_{1}$ and $p_{2}$ (Gorroochurn,, 2018). We will see, by comparing the classical case with the quantum ones, that the requirement of physical indistinguishability has a dramatic impact on the independent behaviour of physical entities that is traditionally assumed at a statistical level.
@@
For the purposes of the present article, we will not dwell on the appearance of Fermi–Dirac statistics in human language (for the interested reader, we refer to (Aerts & Beltran, 2022b, )). We instead intend to clarify the differences between Maxwell–Boltzmann and Bose–Einstein statistics, in particular, with respect to the issue of statistical independence, using a simple but paradigmatic example, taken from human cognition, more specifically, language, as follows (Aerts & Beltran, 2022a, ).
@@ 4 Violation of statistical independence in early quantum theory
Bose–Einstein statistics was discovered before the advent of QM, in a period historiographically known as ‘early quantum theory’ and approximately located between 1900 and 1925. Historians and philosophers of science have carefully reconstructed the intense debate which led to the birth of the new statistics, using both scientific articles and personal correspondence as sources (Klein,, 1961; Howard,, 1990; Darrigol,, 1991; Monaldi,, 2009; Perez & Sauer,, 2010; Gorroochurn,, 2018). In this regard, it is interesting to sketch the different positions of the major players involved in the scientific and epistemological discussions about Bose–Einstein statistics.
@@
The starting point of this story concerned the investigation of how a material entity that is heated begins to emit electromagnetic radiation. An entity that is particularly suited for studies of this type is the ‘black body’, namely, an ideal, but physically realizable with very good approximation, body that absorbs all the electromagnetic radiation incident on it (Eisberg & Resnick,, 1985). In the last years of 1800, it was well known the empirical relationship connecting the energy density and the frequency of the electromagnetic radiation emitted by a black body heated at a given temperature.
@@
Modern manuals of QM typically report that the formal treatment of the interaction between matter and electromagnetic radiation within classical physics leads to a theoretical law, the ‘Raileigh–Jeans law’ which diverges from empirical data at high frequencies, the so-called ‘ultraviolet catastrophe’ (Ehrenfest,, 1911). It was Max Planck who proposed in 1900 what is currently accepted as the correct theoretical law of the black body radiation, in this way determining the birth of early quantum theory (Planck,, 1900). However, the idea of an ultraviolet catastrophe came only after the seminal article of Planck, who was instead inspired by the law of the black body radiation formulated by Wilhelm Wien, which showed good agreement with the empirical data available at the time (Wien,, 1897). Nonetheless, experiments were performed in 1900 which did not agree with Wien’s law in the case of radiation with low frequency (Rubens & Kurlbaum,, 1900).
@@
Looking at a thermodynamic description of the black body radiation, both Wien and Planck were deeply influenced by Boltzmann’s statistical view of thermodynamics (Boltzmann,, 1884). Indeed, Planck derived his law of the black body radiation while trying to provide a thorough derivation of Wien’s law from Boltzmann’s thermodynamics. The introduction of the constant $h$, now named after him, together with the constant $k$, which Planck named after Boltzmann, was primarily intended as a rescue operation for Planck’s thermodynamic theory of the electromagnetic radiation. In an almost miraculous way, Planck’s law, which approximates Wien’s law at high frequencies such that experiments at high frequencies could not show deviations of Wien’s law, also worked very well at low frequencies, thus showing very good agreement with the new empirical data. In modern terms, if we recall the mathematical formulation of Maxwell–Boltzmann and Bose–Einstein statistics introduced in Section 2, we can say that Wien’s law is the Maxwell–Boltzmann approximation of Planck’s law, which instead holds within Bose–Einstein statistics.
@@
The systematic and complete historical account in (Klein,, 1961; Howard,, 1990; Darrigol,, 1991; Monaldi,, 2009; Perez & Sauer,, 2010; Gorroochurn,, 2018) of the scientific debate within early quantum theory reveals that Planck’s idea of ‘quantization’, initially conceived as a mathematical artifice, was still only partial and quite different from the way we understand it nowadays. As a matter of fact, it was Albert Einstein, who applied Planck’s radiation law to the photoelectric effect (Einstein,, 1905), to give a physical meaning to the formula $h\nu$ introduced by Planck and interpret it as the energy of a ‘quantum of light’, later called ‘photon’ (Klein,, 1961).
@@
Both Planck and Einstein solved their respective problems by looking at the way in which ‘particles can be distributed in baskets’ (see Section 3) (Darrigol,, 1991). But, Einstein, and later Paul Ehrenfest, recognised a fundamental problem in the procedure followed by Planck. Indeed, Planck used the method of distributing identical particles into different baskets to find the distribution of energies across the elementary constituents of the electromagnetic radiation. In this way, he found a formula for the black body radiation which coincides with the formula one finds starting directly from the Bose–Einstein distribution. When Einstein studied the photoelectric effect, he attributed a particle nature to those elementary constituents of the radiation, as atoms and molecules. However, it immediately became clear that, if the idea of photons is taken seriously, then the ‘probabilities for the relevant micro-states did not appear to be statistically independent’. As a matter of fact, Ehrenfest later proved rigorously that, assuming statistical independence in a gas of photons, would lead to Wien’s, not Planck’s, law (Ehrenfest,, 1911; Monaldi,, 2009). It is also why Planck, who still relied on Boltzmann’s approach, was not at all inclined to assign a physical reality to these quanta and continued to regard them as abstract elements of a specific thermodynamic view of the situation.
@@
Hence, the positions of Planck and Einstein (together with Ehrenfest) with respect to the interpretation of the radiation law was dramatically divergent. Planck, and with him most other physicists working on related problems, was not troubled by the lack of statistical independence, and thought this was only problematic because Einstein insisted on interpreting these quanta as entities existing in time and space. Einstein and Ehrenfest, on the contrary, considered a serious hindrance the identification of a lack of statistical independence at the level of micro-states, which they could not justify, wanting to interpret these micro-states as physical states of quanta of light, entities existing in time and space, and both looked for an alternative theoretical proposal. This is why, when Einstein learned about Bose’s alternative derivation of Planck’s radiation law, he enthusiastically translated Bose’s article into German and submitted it to Zeischrift für Physik, where the article was published in 1924 (Bose,, 1924). Furthermore, Einstein immediately applied Bose’s approach to an ideal gas of atoms or molecules in a series of articles between 1924 and 1925, which led to the identification of the phenomenon of ‘Bose–Einstein condensation’ (Einstein,, 1924; Einstein, 1925a, ; Einstein, 1925b, ).
@@
This shift of perspective by Einstein also marked the departure from agreement with Ehrenfest, who was not at all enthusiast about Bose’s type of calculation, even being amazed that Einstein embraced it so strongly (see footnote 4) (Perez & Sauer,, 2010; Moskovchenko & Frenkel,, 1990). In the meanwhile, Louis de Broglie had formulated his hypothesis about the wave nature of material entities (Debroglie,, 1924), which Einstein read with great attention. Indeed, in particular in (Einstein, 1925a, ), Einstein saw the statistical dependence of the photons in a gas as caused by a ‘mutual influence’, whose nature was totally unknown but, according to Einstein, de Broglie wave-particle duality could have played a role in it (Monaldi,, 2009; Einstein, 1925a, ). Furthermore, in a letter to Erwin Schrödinger, who was also interested in the investigation of quantum gases (Schrodinger, 1926a, ), Einstein clearly explained that photons do not seem to be independently localized but have a preference to occupy the same micro-state. Equivalently, in Bose–Einstein statistics, the photons crowd together relatively more often than under the hypothesis of statistical independence (Monaldi,, 2009). Hence, Einstein thought about a mysterious force connecting together the photons and causing an unexpected lack of statistical independence.
@@
Einstein embraced Louis de Broglie’s matter waves, albeit with mixed feelings, as they allowed a possible explanation for the lack of statistical independence contained in Planck’s law. Also, the de Broglie’s waves, since they are linked to individual entities, do not stand in the way of the philosophical realism that Einstein prioritized as the criterion for acceptability of an interpretation of the theory. And perhaps, though we suspect this was not at all cleared up for Einstein how that could be possible, they could explain that mysterious lack of statistical independence of photon states. We sense here, in our view, the deep intuition that one also finds in Einstein’s other works, and the courage and stubbornness to stick to it, albeit in opposition to his greatest ally in the struggle for acceptance of the existence of light quanta, Ehrenfest. But Schrödinger’s introduction of waves in the configuration space of the physical entities (Schrodinger, 1926b, ), was for Einstein a bridge too far. Indeed, such waves cannot belong to a reality consistent with Einstein’s philosophical realism, and must trigger deep problems of interpretation in connection with measurement. Einstein was also aware that such waves in configuration space would create a fundamental difficulty for the search of a possible quantum version of his general theory of relativity (Howard,, 1990).
@@
We have seen in Section 2 that modern QM connects Bose–Einstein statistics to the indistinguishability of identical bosons, which is formally represented by the symmetrization postulate. The latter postulate forces an exchange symmetry in the states of the Hilbert space, producing through a mathematical artifice the entangled states that would be responsible of the lack of statistical independence. However, we have also stressed the epistemological difficulties arising from the symmetrization postulate. Moreover, this formal treatment leads to additional difficulties. Indeed, according to the standard interpretation of identical entities in QM, all photons would be ‘completely indistinguishable’, even though they have different energies, e.g., a blue photon would be indistinguishable from a red photon. This assumption does not completely fit with the empirical research on how indistinguishability of photons is used in order to prepare entangled photons and treat them as qubits in quantum computational tasks. What arises in these experiments is that it is sufficient for the photons to be ‘contextually indistinguishable’ when they are measured for them to behave as indistinsguishable bosons (see (Aerts & Beltran,, 2020) and references therein). In other words, photons of different energies are generally considered by quantum experimentalists as distinguishable entities.

### v2

@@ 3 Indistinguishability versus statistical independence
We intend in this section to establish a connection between distinguishability of identical entities and statistical independence of their relevant states. We will first consider the typical situation that occurs in physics [7]. Then, we will see that a completely analogous connection can be made in the case of human cognition, more specifically, language [40, 41].
Traditional derivations of both classical, Maxwell–Boltzmann, and quantum, Bose–Einstein and Fermi–Dirac, distributions employ a method of distributing particles of the same type, which play the role of identical physical entities, across baskets, which play the role of micro-states (see, e.g., [1, 4]). Let us consider a simple situation, namely, the distribution of two physical entities, labelled $1$ and $2$, in two micro-states, labelled $p_{1}$ and $p_{2}$ [7]. We will see, by comparing the classical case with the quantum ones, that the requirement of physical indistinguishability has a dramatic impact on the independent behaviour of physical entities that is traditionally assumed at a statistical level.
@@
For the purposes of the present article, we will not dwell on the appearance of Fermi–Dirac statistics in human language (for the interested reader, we refer to [41]). We instead intend to clarify the differences between Maxwell–Boltzmann and Bose–Einstein statistics, in particular, with respect to the issue of statistical independence, using a simple but paradigmatic example, taken from human cognition, more specifically, language, as follows [40].
@@ 4 Violation of statistical independence in early quantum theory
Bose–Einstein statistics was discovered before the advent of QM, in a period historiographically known as ‘early quantum theory’ and approximately located between 1900 and 1925. Historians and philosophers of science have carefully reconstructed the intense debate which led to the birth of the new statistics, using both scientific articles and personal correspondence as sources [2, 3, 4, 5, 6, 7]. In this regard, it is interesting to sketch the different positions of the major players involved in the scientific and epistemological discussions about Bose–Einstein statistics.
@@
The starting point of this story concerned the investigation of how a material entity that is heated begins to emit electromagnetic radiation. An entity that is particularly suited for studies of this type is the ‘black body’, namely, an ideal, but physically realizable with very good approximation, body that absorbs all the electromagnetic radiation incident on it [48]. In the last years of 1800, it was well known the empirical relationship connecting the energy density and the frequency of the electromagnetic radiation emitted by a black body heated at a given temperature.
@@
Modern manuals of QM typically report that the formal treatment of the interaction between matter and electromagnetic radiation within classical physics leads to a theoretical law, the ‘Raileigh–Jeans law’ which diverges from empirical data at high frequencies, the so-called ‘ultraviolet catastrophe’ [10]. It was Max Planck who proposed in 1900 what is currently accepted as the correct theoretical law of the black body radiation, in this way determining the birth of early quantum theory [9]. However, the idea of an ultraviolet catastrophe came only after the seminal article of Planck, who was instead inspired by the law of the black body radiation formulated by Wilhelm Wien, which showed good agreement with the empirical data available at the time [58]. Nonetheless, experiments were performed in 1900 which did not agree with Wien’s law in the case of radiation with low frequency [59].
@@
Looking at a thermodynamic description of the black body radiation, both Wien and Planck were deeply influenced by Boltzmann’s statistical view of thermodynamics [60]. Indeed, Planck derived his law of the black body radiation while trying to provide a thorough derivation of Wien’s law from Boltzmann’s thermodynamics. The introduction of the constant $h$, now named after him, together with the constant $k$, which Planck named after Boltzmann, was primarily intended as a rescue operation for Planck’s thermodynamic theory of the electromagnetic radiation. In an almost miraculous way, Planck’s law, which approximates Wien’s law at high frequencies such that experiments at high frequencies could not show deviations of Wien’s law, also worked very well at low frequencies, thus showing very good agreement with the new empirical data. In modern terms, if we recall the mathematical formulation of Maxwell–Boltzmann and Bose–Einstein statistics introduced in Section 2, we can say that Wien’s law is the Maxwell–Boltzmann approximation of Planck’s law, which instead holds within Bose–Einstein statistics.
@@
The systematic and complete historical account in [2, 3, 4, 5, 6, 7] of the scientific debate within early quantum theory reveals that Planck’s idea of ‘quantization’, initially conceived as a mathematical artifice, was still only partial and quite different from the way we understand it nowadays. As a matter of fact, it was Albert Einstein, who applied Planck’s radiation law to the photoelectric effect [8], to give a physical meaning to the formula $h\nu$ introduced by Planck and interpret it as the energy of a ‘quantum of light’, later called ‘photon’ [2].
@@
Both Planck and Einstein solved their respective problems by looking at the way in which ‘particles can be distributed in baskets’ (see Section 3) [4]. But, Einstein, and later Paul Ehrenfest, recognised a fundamental problem in the procedure followed by Planck. Indeed, Planck used the method of distributing identical particles into different baskets to find the distribution of energies across the elementary constituents of the electromagnetic radiation. In this way, he found a formula for the black body radiation which coincides with the formula one finds starting directly from the Bose–Einstein distribution. When Einstein studied the photoelectric effect, he attributed a particle nature to those elementary constituents of the radiation, as atoms and molecules. However, it immediately became clear that, if the idea of photons is taken seriously, then the ‘probabilities for the relevant micro-states did not appear to be statistically independent’. As a matter of fact, Ehrenfest later proved rigorously that, assuming statistical independence in a gas of photons, would lead to Wien’s, not Planck’s, law [10, 5]. It is also why Planck, who still relied on Boltzmann’s approach, was not at all inclined to assign a physical reality to these quanta and continued to regard them as abstract elements of a specific thermodynamic view of the situation.
@@
Hence, the positions of Planck and Einstein (together with Ehrenfest) with respect to the interpretation of the radiation law was dramatically divergent. Planck, and with him most other physicists working on related problems, was not troubled by the lack of statistical independence, and thought this was only problematic because Einstein insisted on interpreting these quanta as entities existing in time and space. Einstein and Ehrenfest, on the contrary, considered a serious hindrance the identification of a lack of statistical independence at the level of micro-states, which they could not justify, wanting to interpret these micro-states as physical states of quanta of light, entities existing in time and space, and both looked for an alternative theoretical proposal. This is why, when Einstein learned about Bose’s alternative derivation of Planck’s radiation law, he enthusiastically translated Bose’s article into German and submitted it to Zeischrift für Physik, where the article was published in 1924 [11]. Furthermore, Einstein immediately applied Bose’s approach to an ideal gas of atoms or molecules in a series of articles between 1924 and 1925, which led to the identification of the phenomenon of ‘Bose–Einstein condensation’ [12, 13, 14].
@@
This shift of perspective by Einstein also marked the departure from agreement with Ehrenfest, who was not at all enthusiast about Bose’s type of calculation, even being amazed that Einstein embraced it so strongly (see footnote 4) [6, 46]. In the meanwhile, Louis de Broglie had formulated his hypothesis about the wave nature of material entities [61], which Einstein read with great attention. Indeed, in particular in [13], Einstein saw the statistical dependence of the photons in a gas as caused by a ‘mutual influence’, whose nature was totally unknown but, according to Einstein, de Broglie wave-particle duality could have played a role in it [5, 13].
Now, we have sketched in Section 2 the issues arising from the modern way of framing the description of quantum entities, bosons and fermions, as if this description followed from their complete physical indistinguishability. We stress that these issues started to appear already in the discussions among the founding fathers of QM. Indeed, the experimental realization of Bose–Einstein condensates in which these form spontaneously when the disturbing factor of heat photons is removed by sophisticated techniques of cooling [50, 51], shows that the undisturbed bare state of the gases used is coherent and not optimally random as would follow from both Boltzmann’s and Gibbs’ thermodynamics [62, 63]. Einstein’s articles on gases, with Bose’s calculation, did not go unnoticed in the scientific circles where QM was being pondered but did not get any support. We have mentioned above how negative Ehrenfest’s reaction was to Einstein embracing Bose’s calculation. Einstein, probably looking for support, wrote the following in a letter to Schrödinger, and it is also in this letter he calls Bose’s calculation a ‘statistics’ for the first time: “In Bose’s statistics, which I have used, the quanta or molecules are not treated as independent from one another. […] According to this procedure, the molecules do not seem to be localized independently from each other, but they have a preference to be in the same cell with other molecules. […] According to Bose, the molecules crowd together relatively more often than under the hypothesis of statistical independence” (Einstein to Schrödinger, Berlin, 28 Feb. 1925 [5]).
The new statistics of Bose and Einstein also came to the attention of Planck, who noted that, if the correlations between molecules were confirmed experimentally, this would entail “a fundamental modification of the ordinary conception of the nature and mode of interaction of the molecules” [5, 64]. Schrödinger also responded to the new statistics with an article sharply expressing his disapproval, in the sense that Bose’s procedure represented “a radical departure from the Boltzmann–Gibbs kind of statistics.” Reacting to the explanation he received in Einstein’s letter, he objected that he did not see “for the time being any possibility of understanding the remarkable kind of interaction among the molecules” by which it could be justified [5, 65]. It is also in this article by Schrödinger that the symmetrization postulate (or exchange symmetry principle, see Section 2) was brought forward for the first time with respect to statistics applied to quantum entities. In 1926, and even before he published the article introducing matter waves, Schrödinger wrote a second article directly on Einstein’s gases. Also in this article, Schrödinger remains dismissive of the new Bose–Einstein statistics, which according to him can never be the ‘natural statistics’ of a thermodynamic equilibrium, as it must always be the Boltzmann- or Gibbs-type of statistics with independence of the different micro-states. He also hints in this article that the wave character as put forward in de Broglie’s work might bring solace relative to the situation [5, 15].
Einstein’s stance had not escaped Paul Dirac either. In an unpublished manuscript, he wrote that the molecules were “not distributed independently from one another”, so that there must be “some kind of interaction between them” [5, 66]. In later formulations of the two quantum statistics, Bose–Einstein and Fermi–Dirac, and likewise in the standard introduction that can now be read in most textbooks, the concern about the lack of independence of the micro-states, and the question of what could be an explanation for it, disappeared in its totality. One of the originally highly critical protagonists, Erwin Schrödinger, much later, while working in Ireland having left Austria as a consequence of Nazism, gave a course on the subject of ‘statistical thermodynamics’. A first publication of his course took place in 1944 and a more elaborate book was written in 1952 [67], but Schrödinger’s approach, inspired by the work of Gibbs, is not different from what is now standard in modern textbooks. In these texts, the disappearance of the lack of statistical independence is so profound that theoretical physicists are usually unaware that concerns about it, with the consequent question of what could explain it, played such an important role in the first decades of QM’s development and that almost all of its protagonists reflected explicitly on it.
So, how and when did the turnaround come to the current textbook introduction of quantum statistics, where both bosons and fermions, i.e. ‘all’ quantum entities, are considered to be completely physically indistinguishable? Daniela Monaldi has done historical research with exactly a focus on this question. She mentions as a first root of indistinguishability the awareness of correlations between micro-states when they are stripped of heat-generated randomness, as we have explained above, with Planck’s radiation law as the focus of research and Einstein, Ehrenfest and Planck its main protagonists. As a second root however, she identifies the modification of the classical calculation of the entropy of the monatomic ideal gas, namely, the subtraction of a term depending on the number of possible permutations of identical particles, which is still today frequently justified as a correction required by the symmetry of identical multi-particle entities under their exchange. She also notes that, if in hindsight one can identify these two issues as two roots of indistinguishability, for the purpose of reconstructing the history of quantum statistics these roots are not the plant [5]. Indeed, despite superficial similarities with the modern formulation, the early studies of Planck’s law did not point to the existence of an alternative type of statistics. Furthermore, no one drew any connection between the non-independence of light quanta and an exchange symmetry. This is noteworthy, since exchange symmetry was at that time becoming an issue in ideal gas theory, an issue that is considered in hindsight the second root of indistinguishability.
Monaldi analyzes in detail passages in various articles by scientists who worked in the wake of Einstein, Ehrenfest and Planck on the hypothesis of the existence of light quanta, and where the introduction of a different than the Maxwell–Boltzmann statistics might have been at issue. She concludes that this was not the case. She also notes that the important hypothesis of exchange symmetry has never been discussed in the research of scientists who have worked on light quanta. That the notion of indistinguishability occupies such a prominent place in the standard way of presenting the two quantum statistics is thus primarily connected with the second root, the problem of entropy that does not sum as expected in the situation where gases are mixed. Especially Gibbs in his version of statistical thermodynamics paid much attention to this problem, devoting the entire second chapter of his book to it [5, 63]. Gibbs is also the one who put forward exchange symmetries as important, but his work was not directly related to the emergence of QM taking off with Planck’s work. Nevertheless, because both Planck and Einstein wished to describe gases and light in an integrated way, there was regular influence from prominent figures in gas theory on the group engaged in the development of QM. Even today there is still an overlap of the group of researchers working on what has since been called the ‘Gibbs paradox’ and the typical problems posed by QM (see, e.g., [68]).
We will not elaborate further on this second root in this article since we plan to systematically investigate it in future work, examining to what extent our results related to the Bose–Einstein statistics of human languages are able to shed light on the Gibbs’ paradox. In summary, the technical form of Bose–Einstein and Fermi–Dirac statistics arose as a result of Einstein’s enthusiasm for Bose’s unorthodox method of calculation, but it remained unexplained why this calculation made sense. In the final section of this paper, we will complete our critical look at the modern standard introduction with insights related to our investigation of human language.
@@
Coming back to Einstein, we mentioned that he embraced Louis de Broglie’s matter waves, but with mixed feelings, as they allowed a possible explanation for the lack of statistical independence contained in Planck’s law. Also, the de Broglie’s waves, since they are linked to individual entities, do not stand in the way of the philosophical realism that Einstein prioritized as the criterion for acceptability of an interpretation of the theory. And perhaps, though we suspect this was not at all cleared up for Einstein how that could be possible, they could explain that mysterious lack of statistical independence of photon states. We sense here, in our view, the deep intuition that one also finds in Einstein’s other works, and the courage and stubbornness to stick to it, albeit in opposition to his greatest ally in the struggle for acceptance of the existence of light quanta, Ehrenfest. But Schrödinger’s introduction of waves in the configuration space of the physical entities [69], was for Einstein a bridge too far. Indeed, such waves cannot belong to a reality consistent with Einstein’s philosophical realism, and must trigger deep problems of interpretation in connection with measurement. Einstein was also aware that such waves in configuration space would create a fundamental difficulty for the search of a possible quantum version of his general theory of relativity [3].
@@
We have seen in Section 2 that modern QM connects Bose–Einstein statistics to the indistinguishability of identical bosons, which is formally represented by the symmetrization postulate. The latter postulate forces an exchange symmetry in the states of the Hilbert space, producing through a mathematical artifice the entangled states that would be responsible of the lack of statistical independence. However, we have also stressed the epistemological difficulties arising from the symmetrization postulate. Moreover, this formal treatment leads to additional difficulties. Indeed, according to the standard interpretation of identical entities in QM, all photons would be ‘completely indistinguishable’, even though they have different energies, e.g., a blue photon would be indistinguishable from a red photon. This assumption does not completely fit with the empirical research on how indistinguishability of photons is used in order to prepare entangled photons and treat them as qubits in quantum computational tasks. What arises in these experiments is that it is sufficient for the photons to be ‘contextually indistinguishable’ when they are measured for them to behave as indistinsguishable bosons (see [38] and references therein). In other words, photons of different energies are generally considered by quantum experimentalists as distinguishable entities.